How do I solve for the sides of a rectangle when given only the perimeter (40 feet) and the square feet (90 sq. ft)? Thank you!
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Let represent the width and represent the length.

The perimeter of a rectangle is given by

So we know that which is to say , or, better for our purposes,

The area of a rectangle is the length times the width, so:



Putting the quadratic into standard form:



Solve the quadratic for . Discard the value that produces a negative result when you calculate . Note: This quadratic DOES NOT factor over the rationals. Use the quadratic formula, but leave your answer in simplest radical form.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

How do I solve for the sides of a rectangle when given only the perimeter (40 feet) and the square feet (90 sq. ft)?
-----------------------------------
P = 2(L+W)
40 = 2(L+W)
L + W = 20
L = (20-W)
--------------------
A = LW
90 = (20-W)W
-W^2 +20W = 90
-----
W^2 -20W + 90 = 0
W = 13.16
-----
L = 20-W = 6.84
-----------------------
Cheers,
Stan H.
=========================